OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle. A. 18° B. 36° C. 54° D. 58°

OpenStudy (anonymous):

let one angle be x so another angle =(90-x) now, difference between them is 72 degres so, x-90+x=72 2x=162 x=81, and other angle=(90-81)=9 degrees answer is 81 and 9